OpenStudy (anonymous):

@satellite73 Equation of trajectory

OpenStudy (anonymous):

The equation of a projectile is \[\huge y=\sqrt{3}x - \frac{ gx ^{2} }{ 2 }\] Find the angle of projection. Also find the speed of projection. Where t=o and y=0 also \[\huge \frac{ d ^{2}x }{ dt ^{2} }=0\] and \[\huge \frac{ d ^{2}y }{ dt ^{2} }=0\]

OpenStudy (anonymous):

@Miracrown

OpenStudy (anonymous):

@iambatman

mathslover (mathslover):

Okay. I will try this :-)

OpenStudy (anonymous):

Yes go ahead

mathslover (mathslover):

We need to write down the equation of trajectory first. @No.name can you write it here? We will compare that with the equation given.

OpenStudy (anonymous):

There are 6 forms of theequation of trajectory

mathslover (mathslover):

We will use this form here : \(y = x\tan \theta - \cfrac{gx^2}{2v^2 \cos^2 \theta}\)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

so tan theta is root 3

mathslover (mathslover):

Yeah.. I think so. Do you have any options?

OpenStudy (anonymous):

Yes it is 60

OpenStudy (anonymous):

So finding out speed is easy then

OpenStudy (anonymous):

I didn't get the calculus part

mathslover (mathslover):

Yes, fine. So, we have theta = 60 degrees. Now, \(2v^2 \cos^2 \theta = 2\) \(v^2 \times \cfrac{1}{4} = 1\) \(v^2 = 4\) \(v = \pm 2\)

mathslover (mathslover):

Same here. I'm trying to understand the calculus part.

mathslover (mathslover):

Is v = +/- 2 the answer of the question for speed?

OpenStudy (anonymous):

+2

mathslover (mathslover):

Yeah right!

mathslover (mathslover):

I don't get why the calculus part is given to us here. There is no need for that!

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

Thanks a million, you learnt projectile motion?

mathslover (mathslover):

\(x = vt \cos \theta\) \(\cfrac{dx}{dt} = v\cos theta \times 1\) \(\cfrac{d^2 x}{dt^2 } = v\cos \theta \times 0 \) = 0 Yes, I learnt it earlier.

OpenStudy (anonymous):

Okay, but there are no answers given for the calculus part

mathslover (mathslover):

May be they wanted us to find that ?

OpenStudy (anonymous):

yeah maybe , thanlks!!

mathslover (mathslover):

You're welcome :-)

OpenStudy (loser66):

I don't understand, please explain me from the given equation, it's a downward parabola with the vertex almost (0,0), how can it be a trajectory? if we consider the equation of trajectory, then d^2y / dt^2 = a_y =9.8, it can't be =0 as given information. too many unclear information to me. Please , make it clear. :) @No.name @mathslover

mathslover (mathslover):

@vishweshshrimali5 will help us out.

Parth (parthkohli):

@Loser66 \(vt\cos \theta\) is the x-component in the projectile. The acceleration of the x-component is zero, so mathslover is not wrong.

OpenStudy (loser66):

I am talking about y , not x!! and \(\dfrac{d^2y}{dt^2}=0\) how?

Parth (parthkohli):

Don't worry, I'm getting \(\frac{d^2 y}{dx^2} = g\).

Parth (parthkohli):

Whoops, didn't consider the sign. I meant \(-g\).

OpenStudy (loser66):

One more thing!! I didn't see the work of @mathslover . I was too busy to understand what is wrong with the problem and why we can't use calculus part to solve. I ended up with my questions

mathslover (mathslover):

Wait wait wait.. it is \(\cfrac{d^2y}{\color{blue}{\bf{dt^2}}}\) and NOT \(\cfrac{d^2 y}{\color{blue}{\bf{dx^2}}}\)

mathslover (mathslover):

That's why : \(\cfrac{d^2 y}{dt^2} = 0 \) ...

OpenStudy (loser66):

yes, @mathslover I was taught that for this kind of problem, I have to write 2 equations for x and y. In other words, I have 2 parametric equation for x, y w.r.t time t. x = x_0 + v_(0x)t + 1/2 a_xt^2 , and solve for t to plug back to equation y = y_0 +v_(0y) +1/2a_yt^2 and from the given information d^2x/dt^2 = a_x =0. It's perfectly ok but d^2y/dt^2 =0 is so nonsense

mathslover (mathslover):

\(y = v \sin \theta t - \cfrac{1}{2} gt^2 \) \(\cfrac{dy^2}{dt^2} = -g \) hmm yeah @ParthKohli was right with the RHS.

mathslover (mathslover):

oops d^2y *

Parth (parthkohli):

\[y = vt\sin(\theta) - \dfrac{1}{2}gt^2\]\[y' = v\sin(\theta) - g t\]\[y' ' = -g\]And\[y = x\tan(\theta) - g x^2\dfrac{1}{2v^2 \cos^2 \theta}\]\[y'= \tan(\theta) - g \frac{1}{v^2 \cos^2 \theta } \]\[y'' = 0\]So we basically intermixed our results. lol

OpenStudy (loser66):

it should be that , but he, the Asker, said \(\dfrac{d^2y}{dt^2}=0\) ahaaaah!! he drove me crazy

Parth (parthkohli):

Whoops.

mathslover (mathslover):

both y are different .. one is equation of trajectory and one is for vertical direction.

mathslover (mathslover):

Its getting confusing. Wait, lemme check my book :-)

Parth (parthkohli):

Hmm, both of them are the same. One is a function of \(t\) and another is a function of \(x\), so as to calculate the second derivatives in terms of both.

OpenStudy (loser66):

You see, if we ignore the t, just focus on the given equation, it has no \(y_0\). Does it mean the projectile starts from the ground where y_0 =0, right? but the vertex ( maximum) height is (0,0) hahahaha....

mathslover (mathslover):

I think its wrong..! (the statement) \(\cfrac{d^2 y}{dt^2} \ne 0\) ... as acceleration along y axis can not be sero.

mathslover (mathslover):

*zero

mathslover (mathslover):

\(\cfrac{d^2 x}{dt^2} = 0 \) that is acceleration along x axis is zero. But, if \(\cfrac{d^2 y}{dt^2} = 0 \) then it will conclude that acceleration along x axis and acceleration along y axis is zero. @Loser66 is right. I don't think that the question is right. As, the vertex is zero and the initial position is also zero.

Parth (parthkohli):

Oh.

Parth (parthkohli):

Sorry, I meant the *velocity* is constant. And I read that as d2y/dx2 (so many mistakes argh!). Anyway, we've already shown why d2y/dt2 is -g.

mathslover (mathslover):

Well, IF \(\cfrac{d^2y}{dt^2} = 0\) then \( y = \sqrt{3} x\) which is a straight line.

OpenStudy (loser66):

ok, I satisfy with you guys conclusion. At least, my doubt are off. hihihihi

mathslover (mathslover):

:-)

OpenStudy (anonymous):

The statement is correct , i verified it with my prof.

mathslover (mathslover):

@No.name - Professor verified it? Its impossible bro. See, if you are given with \(\cfrac{d^2y}{dt^2} = 0\) and \(\cfrac{d^2 x }{dt^2 } = -0\) then : For \(\cfrac{d^2y}{dt^2 } = -g = 0 \) and for \(\cfrac{d^2x}{dt^2}= a_x = 0\) Thus, note that -g = 0 or g = 0 The given equation of trajectory becomes : \(y = \sqrt{3} x \) which is an equation of straight line.